### v3

# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability in the Born approximation. The resultant errors are around $20\ \%$ for Ar and $60\ \%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated by the outermost shell electrons. The error for Xe atom will become better for high energy because of the approximation and lack of data.
preprint: KEK-TH-2659
1 Introduction
Dark matter is unknown physics beyond the standard model indicated from astrophysical and cosmological observations (see, e.g., Ref. Murayama:2007ek for review). Dark matter is often assumed to have interactions with the standard model particles. Then, dark matter can be detected by the interaction with probe atoms. The benchmark of dark matter model is called the weakly interacting massive particles (WIMPs) which interacts with nuclei Goodman:1984dc and the abundance is set by the thermal freeze out Lee:1977ua . Direct detection experiments search the WIMP signals from the detection of ionization, scintillation, and heat in the detectors (see Ref. Lewin:1995rx for review).
The Migdal effect migdal1939ionizatsiya ; migdal1941ionizacia ; Landau:1991wop is effective for searching small energy signals from dark matter. The Migdal effect is a inelastic process of dark matter-nucleus scattering and causes the excitation/ionization of electrons from the nuclear recoil. The excitation/ionization of electrons are electromagnetic signals that are easier to detect than heat in the case of elastic scattering. The Migdal effect is not observed in nuclear scattering, but, observed in $\alpha$ decay Rapaport:1975zza ; Rapaport:1975zz ; fischbeck1975spectroscopy ; fischbeck1977angular and $\beta$ decay boehm1954internal ; berlovich1965investigation ; couratin2012first ; Lienard:2015joa ; Fabian:2018pin . Now, experiments for detecting the Migdal effect from nuclear recoils are proceeding Nakamura:2020kex ; Araujo:2022wjh . Recently, the Migdal effect is studied or used in many papers baur1983ionisation ; vegh1983multiple ; wauters1997recoil ; berakdar2001electronic ; Vergados:2005dpd ; Moustakidis:2005gx ; Ejiri:2005aj ; talman2006excitations ; Bernabei:2007jz ; liertzer2012multielectron ; Vergados:2013raa ; pindzola2014neutron ; Sharma:2017fmo ; Ibe:2017yqa ; Dolan:2017xbu ; LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr ; Coskuner:2018are ; EDELWEISS:2019vjv ; Tvrznikova:2019tgx ; CDEX:2019hzn ; XENON:2019zpr ; Baxter:2019pnz ; Essig:2019xkx ; Liang:2019nnx ; Kozaczuk:2020uzb ; SENSEI:2020dpa ; Liu:2020pat ; Nakamura:2020kex ; Pindzola:2020whx ; Kahn:2020fef ; Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu ; LUX:2020yym ; Liang:2020ryg ; He:2020sat ; Flambaum:2020xxo ; Akerib:2021pfd ; Collar:2021fcl ; Bell:2021zkr ; Liao:2021yog ; Menshikov:2021jxa ; Acevedo:2021kly ; COSINE-100:2021poy ; Berlin:2021zbv ; CDEX:2021cll ; Wang:2021oha ; Bell:2021ihi ; Abbamonte:2022rfh ; Chatterjee:2022gbo ; EDELWEISS:2022ktt ; Angevaare:2022zhy ; Campbell-Deem:2022fqm ; Araujo:2022wjh ; DarkSide:2022dhx ; Blanco:2022pkt ; Cox:2022ekg ; GlobalArgonDarkMatter:2022ppc ; Tomar:2022ofh ; Adams:2022zvg ; Berghaus:2022pbu ; Li:2022acp ; NEWS-G:2023qwh ; DarkSide-50:2023fcw ; SuperCDMS:2023sql ; Bell:2023uvf ; Tilly:2023fhw ; AtzoriCorona:2023ais ; Qiao:2023pbw ; Xu:2023wev ; LZ:2023poo ; PandaX:2023xgl ; Yun:2023huf ; Gu:2023pfg ; Li:2023xkf ; Herrera:2023xun ; Kim:2023lkb ; Dzuba:2023zcq ; SENSEI:2023zdf ; Liang:2024tef ; He:2024hkr ; Balan:2024cmq ; MIGDAL:2024alc ; Kang:2024kec ; Ibarra:2024mpq ; Lee:2024wzd , which includes that the comparison LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr to bremsstrahlung of photon Kouvaris:2016afs ; Millar:2018hkv , theoretical comparison to the dark matter-electron scattering Baxter:2019pnz ; Essig:2019xkx , the Migdal-photoabsorption relation Liu:2020pat , formulation for semiconductor response Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu , and the semi-inclusive calculation Cox:2022ekg .
No detection of the Migdal effect from nuclear recoil in liquid Xe probes is reported from Ref. Xu:2023wev . The no detection may caused by the inaccurate predictions for the Migdal transition probability or the signal response in liquid Xe. In this paper, the Migdal probability is compared to the total single-electron inelastic cross section of electron-atom scattering for isolated Ar and Xe. The comparison is possible by expressing the electron-atom scattering cross section by the Migdal probability calculated in Ref. Ibe:2017yqa .
The energy eigenstates of atom are given in Sec. 2. The formalism uses the center of mass coordinates and is useful for treating long-range interactions. The interaction potential is derived in Sec. 3. The interactions are approximately described by the electron-electron interaction. Therefore, the shape of potential is the Coulomb one. The total excitation and ionization cross section is calculated in Sec. 4. Especially, the cross section is expressed by the Migdal transition factor of dark matter-atom scattering for single electron excitation/ionization with the dipole approximation given in Ref. Ibe:2017yqa . Then, the result is compared with a database value which is a combination of experimental and theoretical data. The natural unit is used in this paper.
2 Energy eigenstates of atomic system
The Hamiltonian of atom is
$H_{\rm atom}=\frac{\vec{p}_{n}\,^{2}}{2m_{n}}+\sum_{i=1}^{Z}\frac{\vec{p}_{e,i}\,^{2}}{2m_{e}}-Z\alpha\sum_{i=1}^{Z}\frac{1}{|\vec{r}_{e,i}-\vec{r}_{n}|}+\alpha\sum_{\{i,j\}}\frac{1}{|\vec{r}_{e,i}-\vec{r}_{e,j}|},$ (1)
where $\vec{p}_{n}$ and $\vec{p}_{e,i}$ is the three-dimensional momentum of nucleus and $i$-th electron, $\vec{r}$ is the three-dimensional coordinates, $m_{n}$ and $m_{e}$ are the mass of nucleus and electron, $Z$ is the atomic number, and $\alpha$ is the fine-structure constant.
The above Hamiltonian can be expressed by the center of mass coordinates of atom as
$H_{\rm atom}\approx$ $\frac{\vec{p}_{\rm CM}\,^{2}}{2m_{A}}+\sum_{i=1}^{Z}\frac{\vec{p}_{i}\,^{2}}{2m_{e}}-Z\alpha\sum_{i=1}^{Z}\frac{1}{|\vec{r}_{i}|}+\alpha\sum_{\{i,j\}}\frac{1}{|\vec{r}_{i}-\vec{r}_{j}|}$ (2)
$=$ $\frac{\vec{p}_{\rm CM}\,^{2}}{2m_{A}}+H_{\rm re}$ (3)
with $m_{e}\ll m_{n}$. Here, $\vec{r}_{\rm CM}=(m_{n}\vec{r}_{n}+m_{e}\sum_{i}\vec{r}_{e,i})/(m_{n}+Zm_{e})$ and $\vec{r}_{i}=\vec{r}_{e,i}-\vec{r}_{\rm CM}$ are the center of mass and relative coordinates, $\vec{p}_{\rm CM}\approx m_{A}(\partial_{t}\vec{r}_{\rm CM})$ and $\vec{p}_{i}\approx m_{e}(\partial_{t}\vec{r}_{i})$ are their momentum, and $m_{A}\approx m_{n}+Zm_{e}$ is the atomic mass.
The above Hamiltonian can be separated by the center of mass energy and the others. Therefore, the energy eigenstates of atom are given by $|\Psi_{A}\rangle=|\vec{p}_{\rm CM}\rangle\otimes|\Psi_{\rm re}\rangle$. Here, $|\Psi_{\rm re}\rangle$ is the energy eigenstate of the Hamiltonian for the relative coordinates
$\hat{H}_{\rm re}|\Psi_{\rm re}\rangle=E_{\rm re}|\Psi_{\rm re}\rangle,$ (4)
and $|\vec{p}_{\rm CM}\rangle$ is the plane wave eigenstate for the center of mass state. Note that $H_{\rm re}$ includes the potential of nucleus being at the origin by the approximation. Therefore, the relative coordinates stand for the positions of electrons.
In the following, the energy eigenstate $|\Psi_{\rm re}\rangle$ is assumed to be given by the Slater determinant of single-electron wave function. Also, the single-electron wave functions are expressed by the spherical spinors (or spinor spherical harmonics) (see, e.g., Ref. johnson2007atomic for review).
3 Interaction potential
The Coulomb potential $\phi(\vec{r})$ of atom is described by the Maxwell equation as
$\nabla^{2}\phi(\vec{r})=-\left[Ze\delta(\vec{r}-\vec{r}_{n})-en(\vec{r})\right],$ (1)
where $n(\vec{r})$ is the number density of electrons in an atom and $e$ is the elementary charge. Assuming that the number density is written by the positions of each electron $n(\vec{r})=\sum_{i=1}^{Z}\delta(\vec{r}-\vec{r}_{e,i})$, the solution is
$\phi(\vec{r})=$ $\frac{Ze}{4\pi|\vec{r}-\vec{r}_{n}|}-\sum_{i=1}^{Z}\frac{e}{4\pi|\vec{r}-\vec{r}_{e,i}|}$ (2)
$=$ $\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\left(\frac{Ze}{q^{2}}e^{-i\vec{q}\cdot(\vec{r}-\vec{r}_{n})}-\sum_{i=1}^{Z}\frac{e}{q^{2}}e^{-i\vec{q}\cdot(\vec{r}-\vec{r}_{e,i})}\right).$ (3)
The interaction with the nucleus can be ignored approximately when the excitation and ionization processes are considered. This is because the potential from the nucleus interactions depend on the coordinates as $\vec{r}-\vec{r}_{n}$ and $\vec{r}_{n}$ is related to the relative coordinates $\vec{r}_{i}$ with the suppression factor of $\approx m_{e}/m_{A}$.
Then, the interaction Lagrangian for the ionization scattering of electron and atom is given by
$\mathcal{L}_{\rm int}\approx e\bar{\psi}_{e}\gamma^{\mu}\psi_{e}A_{\mu},$ (4)
where $\psi_{e}$ and $A_{\mu}$ are the electron and photon fields and $\gamma^{\mu}$ is the gamma matrix. The invariant scattering amplitude is given by the $t$-channel and $u$-channel processes as
$\mathcal{M}_{i}\approx 16\pi\alpha m_{e}^{2}\left(\frac{\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\delta_{s_{i}^{F}s_{i}^{I}}}{t_{i}}-\frac{\delta_{s_{i}^{F}s_{\rm in}^{I}}\delta_{s_{\rm in}^{F}s_{i}^{I}}}{u_{i}}\right)$ (5)
under the non-relativistic approximation, where $t_{i}=(p_{{\rm in},I}-p_{{\rm in},F})^{2}$ and $u_{i}=(p_{{\rm in},I}-p_{e,i,F})^{2}$ are the Mandelstam variables for $i$-th electron in the atom. The subscript $I$ and $F$ represent the label for initial and final states, $s$ represents the spin, and an incoming electron is labeled by in.
Then, the invariant amplitude is related to the interaction potential $V(\vec{r}_{e,i}-\vec{r}_{\rm in})$ sakurai1967advanced by the Born approximation. The potential is written as
$V(\vec{r}_{e,i}-\vec{r}_{\rm in})=V_{t}(\vec{r}_{e,i}-\vec{r}_{\rm in})+V_{u}(\vec{r}_{e,i}-\vec{r}_{\rm in}),$ (6)
where Ibe:2017yqa
$V_{t}(\vec{r}_{e,i}-\vec{r}_{\rm in})=$ $-\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}_{t}(q^{2})}{4m_{e}^{2}},$ (7)
$V_{u}(\vec{r}_{e,i}-\vec{r}_{\rm in})=$ $-\int\frac{d^{3}\vec{q}_{u}}{(2\pi)^{3}}e^{i\vec{q}_{u}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}_{u}(q_{u}^{2})}{4m_{e}^{2}}$ (8)
and the invariant scattering amplitude is given as
$\mathcal{M}_{t}(q^{2})=$ $-16\pi\alpha\frac{m_{e}^{2}}{q^{2}},$ (9)
$\mathcal{M}_{u}(q_{u}^{2})=$ $16\pi\alpha\frac{m_{e}^{2}}{q_{u}^{2}}.$ (10)
Here, $s_{i}=(p_{{\rm in},I}+p_{e,i,I})^{2}$ is the Mandelstam variable, $\vec{q}_{u}=\vec{p}_{{\rm in},I}-\vec{p}_{e,i,F}$ is the $u$-channel momentum transfer, $\vec{q}=\vec{p}_{{\rm in},I}-\vec{p}_{{\rm in},F}$ is the momentum transfer of $t$-channel, and $\vec{r}_{\rm in}$ is the coordinates of incoming electron. Remember that the $i$-th electron coordinate $\vec{r}_{e,i}$ is expressed by the relative coordinate $\vec{r}_{i}$ as $\vec{r}_{e,i}=\vec{r}_{i}+\vec{r}_{\rm CM}$.
4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in}\rangle_{I/F}$. The incoming electron states can be taken as plane waves because the atom is regarded to be neutral far from the atom even the atom is ionized. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
$iT_{t,FI}=$ $-2\pi i\delta(E_{F}-E_{I})\sum_{i=1}^{Z}\int d^{3}\vec{r}_{\rm in}d^{3}\vec{r}_{\rm CM}\left(\Pi_{j=1}^{Z}d^{3}\vec{r}_{j}\right)$
×$\times\Phi_{F}^{\dagger}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{k}\})V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})\Phi_{I}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{l}\})$ (1)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\frac{m_{A}}{m_{e}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q}),$ (2)
where the transition matrix of single electron in atom $Z_{FI}(\vec{q})$ is defined as
$Z_{FI}(\vec{q})\equiv$ $\sum_{i=1}^{Z}\int\left(\Pi_{j=1}^{Z}d^{3}\vec{r}_{j}\right)\Psi_{{\rm re},F}^{\dagger}(\{\vec{r}_{k}\})e^{i\vec{q}\cdot\vec{r}_{i}}\Psi_{{\rm re},I}(\{\vec{r}_{l}\})$ (3)
$=$ $\int d^{3}\vec{r}\phi_{E_{e,F}}^{\dagger}(\vec{r})e^{i\vec{q}\cdot\vec{r}}\phi_{E_{e,I}}(\vec{r})$ (4)
$\approx$ $|\vec{q}|z_{FI}(\theta_{qr}).$ (5)
Here, $\phi_{E_{e}}(\vec{r})$ is the spherical spinor of single-electron energy $E_{e}$ and $E_{e}\approx\vec{p}\,^{2}/(2m_{e})$ for the ionized state. The dipole approximation is used in Eq. (5) and the reduced transition matrix element is defined as
$z_{FI}(\theta_{qr})\equiv i\int d^{3}\vec{r}\phi_{E_{e,F}}^{\dagger}(\vec{r})|\vec{r}|\cos\theta_{qr}\phi_{E_{e,I}}(\vec{r}),$ (6)
where $\theta_{qr}$ is the angle between $\vec{q}$ and $\vec{r}$, $\vec{q}=-\vec{p}_{{\rm in},F}+\vec{p}_{{\rm in},I}=\vec{p}_{{\rm CM},F}-\vec{p}_{{\rm CM},I}$ is again the momentum transfer, $\delta^{4}(p_{F}-p_{I})=\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$ shows the four-momentum conservation, and $E_{I/F}$ is the total energy of the initial and final states. Here, this discussion does not use the sudden approximation of Migdal migdal1941ionizacia ; Landau:1991wop , i.e., Galilei transformation of electrons at sudden time, because of the use of center of mass coordinates of atom. Therefore, this calculation is justified even for the case of the Coulomb force. This point is shown for the case of dark matter-nucleus scattering for hydrogen atom in Ref. Kahn:2021ttr .
The $T$-matrix for $u$-channel potential becomes
$iT_{u,FI}=$ $i(2\pi)^{4}\delta(E_{F}-E_{I})\frac{m_{A}}{m_{e}}\int\frac{d^{3}\vec{q}_{u}}{(2\pi)^{3}}\delta^{3}(\vec{q}_{u}+\vec{p}_{{\rm CM},I}-\vec{p}_{{\rm CM},u,F})\mathcal{M}_{u}(q_{u}^{2})$
×$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I}\rangle$
×$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F}|\vec{r}\rangle\phi_{E_{e},I}(\vec{r})e^{i\vec{q}_{u}\cdot\vec{r}}.$ (7)
Here, $\vec{p}_{{\rm CM},u,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ because of the $u$-channel process. The integrations by $\vec{r}_{\rm in}$ and $\vec{r}$ are actually the Fourier transformations of the projected single-electron wave functions. Assuming that the plane wave $\langle\vec{r}|\vec{p}\rangle$ is approximately equal to the Coulomb wave $\phi(\vec{r})$ (see Ref. gozem2015photoelectron for the discussion of the differences), the above equation becomes
$iT_{u,FI}\approx$ $i(2\pi)^{4}\delta(E_{F}-E_{I})\delta^{3}(\vec{p}_{{\rm in},I}+\vec{p}_{{\rm CM},I}-\vec{p}_{{\rm in},F}-\vec{p}_{{\rm CM},F})\delta_{s_{i}^{F}s_{\rm in}^{I}}\frac{m_{A}}{m_{e}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u}).$ (8)
Here, the transition matrix is given as
$Z_{{\rm in}I}(\vec{q}_{u})=\int d^{3}\vec{r}\phi_{{\rm in},F}^{\dagger}(\vec{r})e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}),$ (9)
where $\phi_{{\rm in},F}(\vec{r})$ is the Coulomb wave function with energy for incoming electron and $\vec{q}_{u}=\vec{p}_{{\rm in},I}-\vec{p}_{i,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{{\rm CM},I}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ is again the momentum transfer for $u$-channel process. This approximation will be not good for excitation processes because the excitation states are actually bound states.
Finally, the $T$-matrix is obtained as
$iT_{FI}=$ $iT_{t,FI}+iT_{u,FI}$ (10)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\frac{m_{A}}{m_{e}}\left(\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right).$ (11)
The four-momentum in the laboratory frame is expressed by $p_{{\rm in},I}=(m_{e}+E_{{\rm kin},I},\vec{p}_{{\rm in},I})$, $p_{{\rm in},F}=(m_{e}+E_{{\rm kin},F},\vec{p}_{{\rm in},F})$, $p_{{\rm CM},I}=(m_{A},0)$, and $p_{{\rm CM},F}=(E_{A,F},\vec{p}_{A,F})$, where $E_{\rm kin}\approx\vec{p}_{\rm in}\,^{2}/(2m_{e})$ is the kinetic energy of incoming electron in the laboratory frame. The momentum of final state atom becomes $\vec{p}_{A,F}\approx\vec{p}_{n,F}+\vec{p}_{i,F}$. Here, $\vec{p}_{i,F}=0$ for excitation. Then, $m_{A}=m_{n}+Zm_{e}+E_{{\rm re},I}$ and $E_{A,F}\approx m_{n}+Zm_{e}+E_{{\rm re},F}+\vec{p}_{n,F}\,^{2}/(2m_{n})$ are followed. Finally, the minimum and maximum momentum transfer, $q^{2}_{\rm min}\approx 2m_{e}\Delta E$ and $q^{2}_{\rm max}\approx 2m_{e}(E_{{\rm kin},I}-\Delta E)$, are obtained, where $\Delta E=E_{e,F}-E_{e,I}>0$ is the transition energy of electron in the atom. The $u$-channel momentum transfer becomes $\vec{q}_{u}\,^{2}\approx 2m_{e}E_{{\rm kin},I}-\vec{q}\,^{2}$.
The total inelastic cross section of single-electron scattering is obtained as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{1}{2!}\frac{1}{2}\sum_{F,s_{\rm in}^{F/I}}\int\frac{d^{3}\vec{p}_{{\rm in},F}}{(2\pi)^{3}2p_{{\rm in},F}^{0}}\frac{d^{3}\vec{p}_{{\rm CM},F}}{(2\pi)^{3}2p_{{\rm CM},F}^{0}}\frac{1}{4\sqrt{(p_{{\rm CM},I}\cdot p_{{\rm in},I})^{2}-m_{A}^{2}m_{e}^{2}}}$
×$\times\left(\frac{m_{A}}{m_{e}}\right)^{2}(2\pi)^{4}\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$
×$\times\left|\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right|^{2}.$ (12)
In the cross section, the possible final states $F$ and spins are summed or averaged. The electron spins of final states $s_{i}^{F}$ are included in the label $F$. Note that the sum includes the integration of ionized electron energy. This equation is reduced as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{2\pi\alpha^{2}m_{e}^{2}}{\vec{p}_{{\rm in},I}\,^{2}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{q^{2}_{\rm max}}{q^{2}_{\rm min}}+\left|z_{{\rm in}I}\right|^{2}\log\frac{2m_{e}E_{{\rm kin},I}-q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}-q_{\rm max}^{2}}\right.$
$\left.-2{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{\frac{q_{\rm max}^{2}}{2m_{e}E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}}}\right)\right],$ (13)
or, expressed by the kinetic energy of incoming electron $E_{{\rm kin},I}$ as
$\sigma(E_{{\rm kin},I})=$ $\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{E_{{\rm kin},I}-\Delta E}{\Delta E}\right.$
$\left.-{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{1-\frac{\Delta E}{E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{\Delta E}{E_{{\rm kin},I}}}\right)\right].$ (14)
To obtain Eq. (13), the average/sum of magnetic quantum number for the initial/final states are taken, which makes the factor $\left|z_{FI}(\theta_{qr})\right|^{2}$ isotropic for the angular direction $\left|z_{FI}(\theta_{qr})\right|^{2}\to\left|z_{FI}\right|^{2}$ Ibe:2017yqa . Also, the fact that the transition matrix element does not change the spins is used. In Eq. (14), the relation $\sum_{F}\left|z_{FI}\right|^{2}=\sum_{F}\left|z_{{\rm in}I}\right|^{2}$ is used. The final states of $z_{FI}z_{{\rm in}I}^{*}$ are related by the energy conservation. The first term is actually the cross section by the Coulomb potential without the identification factor of the interacting electrons $1/2!$.
The averaged value of transition probabilities $\sum|z_{FI}|^{2}$ is actually the transition probabilities for the Migdal effect given in Ref. Ibe:2017yqa which uses the Flexible Atomic Code (FAC, cFAC) doi:10.1139/p07-197 . cFAC calculates the probabilities with the Dirac-Hartree-Fock method and a universal central potential.
Figure 1: The total excitation and ionization cross section of electron-atom scattering $\sigma(E_{{\rm kin},I})$ as the function of kinetic energy for incoming electron in laboratory frame $E_{{\rm kin},I}$. The calculation results using the transition probabilities of Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed lines. The red lines show the single excitation case and blue lines show the single ionization case. (top:) For Ar case. The database value for comparison is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The database value for comparison is taken from Morgan Morgan:2024 . The ionization data is limited up to 1 keV.
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the calculation error of ionization is around $20\ \%$ from the database value in 300 eV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around $60\ \%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the error of ionization is around 60 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This error dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term decreases the cross section for low energy and negligible for high energy. Neglected the interference term, the overestimation near the threshold energy is consistent with Ref. ariyasinghe2004total which compares to the Bethe-Born theory inokuti1971inelastic . Note that the probability is dominated by the excitation/ionization from the outermost shell, i.e., $3p$ states for Ar and $5p$ states for Xe.
5 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The errors are $\sim 20\ \%$ for Ar at 10 keV and $\sim 60\ \%$ for Xe compared to the database value shown in Fig. 1. For Xe, the database value is limited and the error will be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. Also, the approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected errors for excitation are not reliable. The differences of threshold energy comes from the error of calculated bound energy. Actually, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ error with experimental data Ibe:2017yqa . Note that this result does not compare the transition probability from the inner shell electrons because the probability is dominated by the outermost electrons.
The electron-nucleus interaction will also cause the excitation/ionization of electrons in small probability. This effect corresponds to the Migdal effect for the Coulomb force, which is neglected in this study because the factor $m_{e}/m_{n}$ suppression is expected in amplitude. Also, there are excitation/ionization processes of $j$-th electron from the interaction with $i$-th electron ($j\neq i$).
An isolated atom is assumed in this calculation, but large volume detectors use the liquid Xe. For electron, the effect of liquid structure becomes effective for $E_{{\rm kin},I}\lesssim 10$ eV and the cross section corresponds to that of gas probes by increasing $E_{{\rm kin},I}$ atrazhev1981hot ; atrazhev1985heating . Then, for high energy electrons $E_{{\rm kin},I}\gg 10$ eV, the cross section will not so differ from the isolated atom calculation. But, the structure of outermost shell may differ between liquid and isolated probes.
Acknowledgements
This work is supported by JSPS Grant-in-Aid for Scientific Research KAKENHI Grant No. JP23KJ2173.