### v6

# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability. The resultant differences are around 30%30\ $\%$ for Ar and 80%80\ $\%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated from the $3p$ shell electrons for Ar. For Xe, $5p$ states dominate the contribution, but, $4d$ states give $\sim 40\ \%$ contribution at 1 keV.
preprint: KEK-TH-2659
1 Introduction
Dark matter is unknown physics beyond the standard model indicated from astrophysical and cosmological observations (see, e.g., Ref. Murayama:2007ek for review). Dark matter is often assumed to have interactions with the standard model particles. Then, dark matter can be detected by the interaction with probe atoms. The benchmark of dark matter model is called the weakly interacting massive particles (WIMPs) which interacts with nuclei Goodman:1984dc and the abundance is set by the thermal freeze out Lee:1977ua . Direct detection experiments search the WIMP signals from the detection of ionization, scintillation, and heat in the detectors (see Ref. Lewin:1995rx for review).
The Migdal effect migdal1939ionizatsiya ; migdal1941ionizacia ; Landau:1991wop is an inelastic process of scattered nucleus. It causes the excitation/ionization of electrons from the nuclear recoil or decay. The Migdal effect is observed in $\alpha$ decay Rapaport:1975zza ; Rapaport:1975zz ; fischbeck1975spectroscopy ; fischbeck1977angular and $\beta$ decay boehm1954internal ; berlovich1965investigation ; couratin2012first ; Lienard:2015joa ; Fabian:2018pin , however, not observed in nuclear scattering. Now, experiments for detecting the Migdal effect from gas probes are proceeding Nakamura:2020kex ; Araujo:2022wjh .
The Migdal effect is also effective for searching small energy signals from dark matter. The Migdal effect occurs as an inelastic process of dark matter-nucleus scattering and causes the excitation/ionization of electrons from the nuclear recoil. The excitation/ionization of electrons returns electromagnetic signals that are easier to detect than heat in the case of elastic scattering. Recently, the Migdal effect is studied or used in many papers baur1983ionisation ; vegh1983multiple ; wauters1997recoil ; berakdar2001electronic ; Vergados:2005dpd ; Moustakidis:2005gx ; Ejiri:2005aj ; talman2006excitations ; Bernabei:2007jz ; liertzer2012multielectron ; Vergados:2013raa ; pindzola2014neutron ; Sharma:2017fmo ; Ibe:2017yqa ; Dolan:2017xbu ; LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr ; Coskuner:2018are ; EDELWEISS:2019vjv ; Tvrznikova:2019tgx ; CDEX:2019hzn ; XENON:2019zpr ; Baxter:2019pnz ; Essig:2019xkx ; Liang:2019nnx ; Kozaczuk:2020uzb ; SENSEI:2020dpa ; Liu:2020pat ; Nakamura:2020kex ; Pindzola:2020whx ; Kahn:2020fef ; Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu ; LUX:2020yym ; Liang:2020ryg ; He:2020sat ; Flambaum:2020xxo ; Akerib:2021pfd ; Collar:2021fcl ; Bell:2021zkr ; Liao:2021yog ; Menshikov:2021jxa ; Acevedo:2021kly ; COSINE-100:2021poy ; Berlin:2021zbv ; CDEX:2021cll ; Wang:2021oha ; Bell:2021ihi ; Abbamonte:2022rfh ; Chatterjee:2022gbo ; EDELWEISS:2022ktt ; Angevaare:2022zhy ; Campbell-Deem:2022fqm ; Araujo:2022wjh ; DarkSide:2022dhx ; Blanco:2022pkt ; Cox:2022ekg ; GlobalArgonDarkMatter:2022ppc ; Tomar:2022ofh ; Adams:2022zvg ; Berghaus:2022pbu ; Li:2022acp ; NEWS-G:2023qwh ; DarkSide-50:2023fcw ; SuperCDMS:2023sql ; Bell:2023uvf ; Tilly:2023fhw ; AtzoriCorona:2023ais ; Qiao:2023pbw ; Xu:2023wev ; LZ:2023poo ; PandaX:2023xgl ; Yun:2023huf ; Gu:2023pfg ; Li:2023xkf ; Herrera:2023xun ; Kim:2023lkb ; Dzuba:2023zcq ; SENSEI:2023zdf ; Liang:2024tef ; He:2024hkr ; Balan:2024cmq ; MIGDAL:2024alc ; Kang:2024kec ; Ibarra:2024mpq ; Lee:2024wzd , which includes that the comparison LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr to bremsstrahlung of photon Kouvaris:2016afs ; Millar:2018hkv , theoretical comparison to the dark matter-electron scattering Baxter:2019pnz ; Essig:2019xkx , the Migdal-photoabsorption relation Liu:2020pat , formulation for semiconductor response Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu , and the semi-inclusive calculation Cox:2022ekg .
No detection of the Migdal effect from nuclear recoil in liquid Xe probes is reported from Ref. Xu:2023wev . The no detection may caused by the inaccurate predictions for the Migdal transition probability or the signal response in liquid Xe. In this paper, the Migdal probability is compared to the total single-electron inelastic cross section of electron-atom scattering for isolated Ar and Xe. The comparison is possible by expressing the electron-atom scattering cross section by the Migdal probability calculated in Ref. Ibe:2017yqa .
The energy eigenstates of atom are given in Sec. 2. The formalism uses the center of mass coordinates and is useful for treating long-range interactions. The interaction potential is derived in Sec. 3. The interactions are approximately described by the electron-electron interaction. Therefore, the shape of potential is the Coulomb one. The total excitation and ionization cross section is calculated in Sec. 4. Especially, the cross section is expressed by the Migdal transition factor of dark matter-atom scattering for single electron excitation/ionization with the dipole approximation given in Ref. Ibe:2017yqa . Then, the result is compared with a database value which is a combination of experimental and theoretical data and other theoretical calculations. The natural unit is used in this paper.
2 Energy eigenstates of atomic system
The Hamiltonian of atom is
$H_{\rm atom}=\frac{\vec{p}_{n}\,^{2}}{2m_{n}}+\sum_{i=1}^{Z}\frac{\vec{p}_{e,i}\,^{2}}{2m_{e}}-Z\alpha\sum_{i=1}^{Z}\frac{1}{|\vec{r}_{e,i}-\vec{r}_{n}|}+\alpha\sum_{\{i,j\}}\frac{1}{|\vec{r}_{e,i}-\vec{r}_{e,j}|},$ (1)
where $\vec{p}_{n}$ and $\vec{p}_{e,i}$ is the three-dimensional momentum of the nucleus and $i$-th electron, $\vec{r}$ is the three-dimensional coordinates, $m_{n}$ and $m_{e}$ are the mass of the nucleus and electron, $Z$ is the atomic number, and $\alpha$ is the fine structure constant. The bracket $\{i,j\}$ represents the sum of the pair $i$ and $j$ under $i<j$.
The above Hamiltonian can be expressed by the center of mass coordinates of atom as
$H_{\rm atom}\approx$ $\frac{\vec{p}_{\rm CM}\,^{2}}{2m_{A}}+\sum_{i=1}^{Z}\frac{\vec{p}_{i}\,^{2}}{2m_{e}}-Z\alpha\sum_{i=1}^{Z}\frac{1}{|\vec{r}_{i}|}+\alpha\sum_{\{i,j\}}\frac{1}{|\vec{r}_{i}-\vec{r}_{j}|}$ (2)
$=$ $\frac{\vec{p}_{\rm CM}\,^{2}}{2m_{A}}+H_{\rm re}$ (3)
with $m_{e}\ll m_{n}$. Here, $\vec{r}_{\rm CM}=(m_{n}\vec{r}_{n}+m_{e}\sum_{i}\vec{r}_{e,i})/(m_{n}+Zm_{e})$ and $\vec{r}_{i}=\vec{r}_{e,i}-\vec{r}_{\rm CM}$ are the center of mass and relative coordinates, $\vec{p}_{\rm CM}\approx m_{A}(\partial_{t}\vec{r}_{\rm CM})$ and $\vec{p}_{i}\approx m_{e}(\partial_{t}\vec{r}_{i})$ are their momentum, and $m_{A}\approx m_{n}+Zm_{e}$ is the atomic mass.
The above Hamiltonian can be separated by the center of mass energy and the others. Therefore, the energy eigenstates of the atom are given by $|\Psi_{A}\rangle=|\vec{p}_{\rm CM}\rangle\otimes|\Psi_{\rm re}\rangle$. Here, $|\Psi_{\rm re}\rangle$ is the energy eigenstate of the Hamiltonian for the relative coordinates
$\hat{H}_{\rm re}|\Psi_{\rm re}\rangle=E_{\rm re}|\Psi_{\rm re}\rangle,$ (4)
and $|\vec{p}_{\rm CM}\rangle$ is the plane wave eigenstate for the center of mass state. Note that $H_{\rm re}$ includes the potential of nucleus being at the origin by approximation. Therefore, the relative coordinates represent the positions of electrons.
In the following, the energy eigenstate $|\Psi_{\rm re}\rangle$ is assumed to be given by the single Slater determinant of the single-electron wave function. Also, the single-electron wave functions are expressed by the spherical spinors (or spinor spherical harmonics) (see, e.g., Ref. johnson2007atomic for review).
3 Interaction potential
The Coulomb potential $\phi(\vec{r})$ of atom is described by the Maxwell equation as
$\nabla^{2}\phi(\vec{r})=-\left[Ze\delta(\vec{r}-\vec{r}_{n})-en(\vec{r})\right],$ (1)
where $n(\vec{r})$ is the number density of electrons in an atom and $e$ is the elementary charge. Assuming that the number density is written by the positions of each electron $n(\vec{r})=\sum_{i=1}^{Z}\delta(\vec{r}-\vec{r}_{e,i})$, the solution is
$\phi(\vec{r})=$ $\frac{Ze}{4\pi|\vec{r}-\vec{r}_{n}|}-\sum_{i=1}^{Z}\frac{e}{4\pi|\vec{r}-\vec{r}_{e,i}|}$ (2)
$=$ $\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\left(\frac{Ze}{q^{2}}e^{-i\vec{q}\cdot(\vec{r}-\vec{r}_{n})}-\sum_{i=1}^{Z}\frac{e}{q^{2}}e^{-i\vec{q}\cdot(\vec{r}-\vec{r}_{e,i})}\right).$ (3)
The interaction with the nucleus can be ignored approximately when the excitation and ionization processes are considered. This is because the potential from the nucleus interactions depend on the coordinates as $\vec{r}-\vec{r}_{n}$ and $\vec{r}_{n}$ is related to relative coordinates $\vec{r}_{i}$ with the suppression factor of $\approx m_{e}/m_{n}$.
Then, the interaction Lagrangian for the ionization scattering of electron and atom is given by
$\mathcal{L}_{\rm int}\approx e\bar{\psi}_{e}\gamma^{\mu}\psi_{e}A_{\mu},$ (4)
where $\psi_{e}$ and $A_{\mu}$ are the electron and photon fields and $\gamma^{\mu}$ is the gamma matrix. The invariant scattering amplitude is given by the $t$-channel and $u$-channel processes as
$\mathcal{M}_{i}\approx 16\pi\alpha m_{e}^{2}\left(\frac{\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\delta_{s_{i}^{F}s_{i}^{I}}}{t_{i}}-\frac{\delta_{s_{i}^{F}s_{\rm in}^{I}}\delta_{s_{\rm in}^{F}s_{i}^{I}}}{u_{i}}\right)$ (5)
under the non-relativistic approximation, where $t_{i}=(p_{{\rm in},I}-p_{{\rm in},F})^{2}$ and $u_{i}=(p_{{\rm in},I}-p_{e,i,F})^{2}$ are the Mandelstam variables for $i$-th electron in the atom. The subscript $I$ and $F$ represent the label for initial and final states, $s$ represents the spin, and an incoming electron is labeled by in.
Then, the invariant amplitude is related to the interaction potential $V(\vec{r}_{e,i}-\vec{r}_{\rm in})$ sakurai1967advanced by the Born approximation. The potential is written as
$V(\vec{r}_{e,i}-\vec{r}_{\rm in})=V_{t}(\vec{r}_{e,i}-\vec{r}_{\rm in})+V_{u}(\vec{r}_{e,i}-\vec{r}_{\rm in}),$ (6)
where Ibe:2017yqa
$V_{t}(\vec{r}_{e,i}-\vec{r}_{\rm in})=$ $-\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}_{t}(q^{2})}{4m_{e}^{2}},$ (7)
$V_{u}(\vec{r}_{e,i}-\vec{r}_{\rm in})=$ $-\int\frac{d^{3}\vec{q}_{u}}{(2\pi)^{3}}e^{i\vec{q}_{u}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}_{u}(q_{u}^{2})}{4m_{e}^{2}}$ (8)
and the invariant scattering amplitude is given as
$\mathcal{M}_{t}(q^{2})=$ $-16\pi\alpha\frac{m_{e}^{2}}{q^{2}},$ (9)
$\mathcal{M}_{u}(q_{u}^{2})=$ $16\pi\alpha\frac{m_{e}^{2}}{q_{u}^{2}}.$ (10)
Here, $s_{i}=(p_{{\rm in},I}+p_{e,i,I})^{2}$ is the Mandelstam variable, $\vec{q}_{u}=\vec{p}_{{\rm in},I}-\vec{p}_{e,i,F}$ is the $u$-channel momentum transfer, $\vec{q}=\vec{p}_{{\rm in},I}-\vec{p}_{{\rm in},F}$ is the momentum transfer of $t$-channel, and $\vec{r}_{\rm in}$ is the coordinates of incoming electron. Remember that the $i$-th electron coordinate $\vec{r}_{e,i}$ is expressed by the relative coordinate $\vec{r}_{i}$ as $\vec{r}_{e,i}=\vec{r}_{i}+\vec{r}_{\rm CM}$.
The above assumptions of a single Slater determinant and pure Coulomb potential neglect the electron correlation. These assumptions are not good in low energy scattering region and better in high energy region peixoto1969elastic ; naon1972importance ; khatun2021theoretical ; parvin2023theoretical . This is because the incident electron passes through the atom in a short time. Ref. peixoto1969elastic shows that those approximations look good for $40$ keV of incident energy and the discrepancy is about 6-12 % in small scattering angle region and better in large angle region for He and Ne atoms. In addition, this approach partially incorporates electron screening and correlation effects beyond the pure Coulomb potential and goes beyond a hydrogenic approximation.
4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in},s_{\rm in}\rangle_{I/F}$. Incoming electron states are taken as plane waves. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
$iT_{t,FI}=$ $-2\pi i\delta(E_{F}-E_{I})\sum_{i=1}^{Z}\int d^{3}\vec{r}_{\rm in}d^{3}\vec{r}_{\rm CM}\left(\Pi_{j=1}^{Z}d^{3}\vec{r}_{j}\right)$
$\times\Phi_{F}^{\dagger}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{k}\})V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})\Phi_{I}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{l}\})$ (1)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\frac{m_{A}}{m_{e}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q}),$ (2)
where the transition matrix of single electron in atom $Z_{FI}(\vec{q})$ is defined as
$Z_{FI}(\vec{q})\equiv$ $\sum_{i=1}^{Z}\int\left(\Pi_{j=1}^{Z}d^{3}\vec{r}_{j}\right)\Psi_{{\rm re},F}^{\dagger}(\{\vec{r}_{k}\})e^{i\vec{q}\cdot\vec{r}_{i}}\Psi_{{\rm re},I}(\{\vec{r}_{l}\})$ (3)
$=$ $\int d^{3}\vec{r}\phi_{E_{e,F}}^{\dagger}(\vec{r})e^{i\vec{q}\cdot\vec{r}}\phi_{E_{e,I}}(\vec{r})$ (4)
$\approx$ $|\vec{q}|z_{FI}(\theta_{qr}).$ (5)
Here, $\phi_{E_{e}}(\vec{r})$ is the spherical spinor of single-electron energy $E_{e}$ and $E_{e}\approx\vec{p}\,^{2}/(2m_{e})$ for the ionized state. The dipole approximation is used in Eq. (5) and the reduced transition matrix element is defined as
$z_{FI}(\theta_{qr})\equiv i\int d^{3}\vec{r}\phi_{E_{e,F}}^{\dagger}(\vec{r})|\vec{r}|\cos\theta_{qr}\phi_{E_{e,I}}(\vec{r}),$ (6)
where $\theta_{qr}$ is the angle between $\vec{q}$ and $\vec{r}$, $\vec{q}=-\vec{p}_{{\rm in},F}+\vec{p}_{{\rm in},I}=\vec{p}_{{\rm CM},F}-\vec{p}_{{\rm CM},I}$ is again the momentum transfer, $\delta^{4}(p_{F}-p_{I})=\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$ shows the four-momentum conservation, and $E_{I/F}$ is the total energy of the initial and final states. Here, this discussion does not use the sudden approximation of Migdal migdal1941ionizacia ; Landau:1991wop , i.e., Galilei transformation of electrons at sudden time, because of the use of center of mass coordinates of atom. Therefore, this calculation is justified even for the case of the Coulomb force. This point is shown for the case of dark matter-nucleus scattering for hydrogen atom in Ref. Kahn:2021ttr .
The $T$-matrix for $u$-channel potential becomes
$iT_{u,FI}=$ $i(2\pi)^{4}\delta(E_{F}-E_{I})\frac{m_{A}}{m_{e}}\int\frac{d^{3}\vec{q}_{u}}{(2\pi)^{3}}\delta^{3}(\vec{q}_{u}+\vec{p}_{{\rm CM},I}-\vec{p}_{{\rm CM},u,F})\mathcal{M}_{u}(q_{u}^{2})$
$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I},s_{{\rm in},I}\rangle$
$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F},s_{{\rm in},F}|\vec{r}\rangle e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}).$ (7)
Here, $\vec{p}_{{\rm CM},u,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ because of the $u$-channel process. The integrations by $\vec{r}_{\rm in}$ and $\vec{r}$ are actually the Fourier transformations of the projected single-electron wave functions. Assuming that the plane wave $\langle\vec{r}\,|\vec{p},s\rangle$ is approximately equal to the Coulomb wave $\phi(\vec{r})$ (see Ref. gozem2015photoelectron for the discussion of the differences), the above equation becomes
$iT_{u,FI}\approx$ $i(2\pi)^{4}\delta(E_{F}-E_{I})\delta^{3}(\vec{p}_{{\rm in},I}+\vec{p}_{{\rm CM},I}-\vec{p}_{{\rm in},F}-\vec{p}_{{\rm CM},F})\delta_{s_{i}^{F}s_{\rm in}^{I}}\frac{m_{A}}{m_{e}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u}).$ (8)
Here, the transition matrix is given as
$Z_{{\rm in}I}(\vec{q}_{u})=\int d^{3}\vec{r}\phi_{{\rm in},F}^{\dagger}(\vec{r})e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}),$ (9)
where $\phi_{{\rm in},F}(\vec{r})$ is the Coulomb wave function with energy for incoming electron and $\vec{q}_{u}=\vec{p}_{{\rm in},I}-\vec{p}_{i,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{{\rm CM},I}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ is again the momentum transfer for $u$-channel process. This approximation will be not good for excitation processes because the excitation states are actually bound states.
Finally, the $T$-matrix is obtained as
$iT_{FI}=$ $iT_{t,FI}+iT_{u,FI}$ (10)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\frac{m_{A}}{m_{e}}\left(\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right).$ (11)
The four-momentum in the laboratory frame is expressed by $p_{{\rm in},I}=(m_{e}+E_{{\rm kin},I},\vec{p}_{{\rm in},I})$, $p_{{\rm in},F}=(m_{e}+E_{{\rm kin},F},\vec{p}_{{\rm in},F})$, $p_{{\rm CM},I}=(m_{A},0)$, and $p_{{\rm CM},F}=(E_{A,F},\vec{p}_{A,F})$, where $E_{\rm kin}\approx\vec{p}_{\rm in}\,^{2}/(2m_{e})$ is the kinetic energy of incoming electron in the laboratory frame. The momentum of final state atom becomes $\vec{p}_{A,F}\approx\vec{p}_{n,F}+\vec{p}_{i,F}$. Here, $\vec{p}_{i,F}=0$ for excitation. Then, $m_{A}=m_{n}+Zm_{e}+E_{{\rm re},I}$ and $E_{A,F}\approx m_{n}+Zm_{e}+E_{{\rm re},F}+\vec{p}_{n,F}\,^{2}/(2m_{n})$ are followed. Finally, the minimum and maximum momentum transfer, $q^{2}_{\rm min}\approx 2m_{e}\Delta E$ and $q^{2}_{\rm max}\approx 2m_{e}(E_{{\rm kin},I}-\Delta E)$, are obtained, where $\Delta E=E_{e,F}-E_{e,I}>0$ is the transition energy of electron in the atom. The $u$-channel momentum transfer becomes $\vec{q}_{u}\,^{2}\approx 2m_{e}E_{{\rm kin},I}-\vec{q}\,^{2}$.
The total inelastic cross section of single-electron scattering is obtained as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{1}{2!}\frac{1}{2}\sum_{F,s_{\rm in}^{F/I}}\int\frac{d^{3}\vec{p}_{{\rm in},F}}{(2\pi)^{3}2p_{{\rm in},F}^{0}}\frac{d^{3}\vec{p}_{{\rm CM},F}}{(2\pi)^{3}2p_{{\rm CM},F}^{0}}\frac{1}{4\sqrt{(p_{{\rm CM},I}\cdot p_{{\rm in},I})^{2}-m_{A}^{2}m_{e}^{2}}}$
$\times\left(\frac{m_{A}}{m_{e}}\right)^{2}(2\pi)^{4}\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$
$\times\left|\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right|^{2}.$ (12)
In the cross section, the possible final states $F$ and spins are summed or averaged. The electron spins of final states $s_{i}^{F}$ are included in the label $F$. The sum of the initial spin for atomic electrons are included in $Z_{FI}$ Ibe:2017yqa . Note that the sum includes the integration of ionized electron energy. This equation is reduced as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{4\pi\alpha^{2}m_{e}^{2}}{\vec{p}_{{\rm in},I}\,^{2}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{q^{2}_{\rm max}}{q^{2}_{\rm min}}+\left|z_{{\rm in}I}\right|^{2}\log\frac{2m_{e}E_{{\rm kin},I}-q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}-q_{\rm max}^{2}}\right$.
$\left.-2{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{\frac{q_{\rm max}^{2}}{2m_{e}E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}}}\right)\right],$ (13)
or, expressed by the kinetic energy of incoming electron $E_{{\rm kin},I}$ as
$\sigma(E_{{\rm kin},I})=$ $\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left[\left(\left|z_{FI}\right|^{2}+\left|z_{{\rm in}I}\right|^{2}\right)\log\frac{E_{{\rm kin},I}-\Delta E}{\Delta E}\right$.
$\left.-2{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{1-\frac{\Delta E}{E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{\Delta E}{E_{{\rm kin},I}}}\right)\right].$ (14)
To obtain Eq. (13), the average/sum of magnetic quantum number for the initial/final states are taken, which makes the factor $\left|z_{FI}(\theta_{qr})\right|^{2}$ isotropic for the angular direction $\left|z_{FI}(\theta_{qr})\right|^{2}\to\left|z_{FI}\right|^{2}$ Ibe:2017yqa . Also, the fact that the transition matrix element does not change the spins is used. The final states of $z_{FI}z_{{\rm in}I}^{*}$ are related by the energy conservation as $E_{{\rm in},F}\approx E_{{\rm kin},I}-\Delta E$. The log term shows the kinematical constraint from the phase space of $E_{e,F}$ as $E_{{\rm kin},I}-2\Delta E>0$. The $\left|z_{FI}\right|^{2}$ part in the first term is actually the scattering cross section by the Coulomb potential without the identification factor of the interacting electrons $1/2!$.
The averaged value of $|z_{FI}|^{2}$ is actually the transition probabilities for the Migdal effect given in Ref. Ibe:2017yqa which uses the Flexible Atomic Code (FAC, cFAC) doi:10.1139/p07-197 for the calculation. cFAC calculates the probabilities with the Dirac-Hartree-Fock method and a universal central potential.
Figure 1: The total excitation and ionization cross section of electron-atom scattering $\sigma(E_{{\rm kin},I})$ as the function of kinetic energy for incoming electron in laboratory frame $E_{{\rm kin},I}$. The calculation results using the transition probabilities of Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed lines. The red lines show the single excitation case and blue lines show the single ionization case. (top:) For Ar case. The database value for comparison is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The database value for comparison is taken from Morgan Morgan:2024 . The ionization data is limited up to 1 keV.
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the difference of ionization cross section is around 30%30\ $\%$ from the database value in 5 keV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around 60%60\ $\%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the difference of ionization cross section is around 80 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This gap dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term is effective at low energy and negligible at high energy. Note that the probability is almost dominated by the excitation/ionization from the outermost shell. For Ar, $3p$ states dominate the contribution and $2p$ and $3s$ states give $\sim 1\ \%$ contribution at $E_{{\rm kin},I}=10$ keV. For Xe, $5p$ states dominate up to $\sim$ 100 eV, but, $4d$ states have $\sim 40\ \%$ contribution at $E_{{\rm kin},I}=1$ keV.
5 Comparison with other results
The inelastic cross section is also calculated by using the Bethe theory inokuti1971inelastic ; inokuti1975total ; inokuti1981oscillator ; salvat2022inelastic . The obtained formula of inelastic cross section for Ar is inokuti1975total
$\sigma_{\rm Inokuti}(E_{{\rm kin},I})=4\pi a_{0}^{2}\frac{\rm Ryd}{E_{{\rm kin},I}}\left[M_{\rm Inokuti}^{2}\log\left(4\frac{E_{{\rm kin},I}}{\rm Ryd}\right)+M_{\rm Inokuti}^{2}\log c_{\rm Inokuti}+\gamma_{\rm Inokuti}\frac{\rm Ryd}{E_{{\rm kin},I}}\right],$ (1)
where $a_{0}=1/(m_{e}\alpha)$ is the Bohr radius, Ryd = 13.605 eV is the Rydberg energy, $\gamma_{\rm Inokuti}=Z(-7/4+\log(B/E_{{\rm kin},I}))$ for electron, and $B$ is the average binding energy which is taken as 13 eV because the 3p states dominate the signal. $M_{\rm Inokuti}^{2}=5.4755$ and $M_{\rm Inokuti}^{2}\log c_{\rm Inokuti}=-1.063$ are taken from the Hartree-Fock calculation (see Ref. inokuti1975total ). Ref. inokuti1975total ; inokuti1981oscillator calculated the ionization cross section from hydrogen ($Z=1$) to strontium ($Z=38$).
For Xe, the result of Ref. salvat2022inelastic is used. It shows
$\sigma_{\rm Salvat}(\beta^{2})=\frac{2\pi\alpha^{2}}{m_{e}\beta^{2}}\left[M_{\rm Salvat}^{2}\left(\log\frac{\beta^{2}}{1-\beta^{2}}-\beta^{2}\right)+C_{\rm Salvat}\right],$ (2)
where $\beta$ is the velocity of incoming electron and the values of $M_{\rm Salvat}^{2}$ and $C_{\rm Salvat}$ are taken from the figure in Ref. salvat2022inelastic . This calculation used the shell corrections, self-consistent Dirac-Hartree-Fock-Slater potential, independent-electron approximation, and relativistic plane-wave Born approximation. They reported less than 1 % precision for $E\gtrsim 100$ MeV and $\sim 8$ % precision for $E\sim 5$ MeV for high $Z$ atoms and good agreement with the results of Ref. inokuti1975total ; inokuti1981oscillator .
Figure 2: The comparison of total inelastic cross section $\sigma(E_{{\rm kin},I})$. The calculation results using Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed and dotted lines. (top:) For Ar case. The dotted lines are from the theoretical calculation in Ref. inokuti1975total with the Hartree-Fock calculation. The database value is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The dotted lines are from the theoretical calculation in Ref. salvat2022inelastic and not convergent at low energy region because of its formula. The database value is taken from Morgan Morgan:2024 .
The comparison result is shown in Fig. 2. For both comparisons, this result shows smaller value than the database value and Ref. inokuti1975total ; salvat2022inelastic . Also, the result from Ref. salvat2022inelastic gets consistent parameters with that from Ref. inokuti1975total . The line of Ref. salvat2022inelastic aims over MeV region with heavy incoming particles and does not converge at low energy region because of its shape of formula. Their result will be applied here because electrons are not fast in keV.
The binding energies and wave functions of electrons are calculated with electron correlations, e.g., Ref. Baxter:2019pnz shows $\sim 5$ % precision for binding energies which is better than Ref. Ibe:2017yqa . Also, Ref. Cox:2022ekg shows $\sim 20$ % differences of ionization probabilities for xenon 5s and argon 3p states with a 20 keV ionized electron. These uncertainties are significant for outer shell states and this comparison method actually relies on the precision of outermost shell states. Note that the uncertainty becomes much smaller for inner shell electrons. Density functional theory may provide better precision of binding energies, but, it does not provide the accurate wave functions.
In addition, this scattering formulation approximates the electron correlations, but this uncertainty is difficult to estimate. The use of $u$-channel potential will be valid for $E_{{\rm kin},I}\gtrsim 1$ keV electron because faster electrons feel less effect of atomic electron clouds than slower ones.
6 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The differences are $\sim 30\ \%$ for Ar at 10 keV and $\sim 80\ \%$ for Xe compared with the database value shown in Fig. 1. For Xe, the database value is limited and the differences may be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. This approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected differences for excitation are not reliable. The differences in the low energy threshold come from the kinematical constraint, $E_{{\rm kin},I}>2\Delta E$. In addition, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ differences from the experimental data Ibe:2017yqa . In this calculation, the change of the electron mass in $m_{A}$ by scattering is neglected. This change will cause a difference of $m_{e}/m_{n}\sim 10^{-4}$.
The electron-nucleus interaction will also cause the excitation/ionization of electrons in small probability. This effect corresponds to the Migdal effect for the Coulomb force, which is neglected in this study because the factor $m_{e}/m_{n}$ suppression is expected in amplitude. Also, there are excitation/ionization processes of $j$-th electron from the interaction with $i$-th electron ($j\neq i$).
An isolated atom is assumed in this calculation, but large volume detectors use the liquid Xe. For electron, the effect of liquid structure becomes effective for $E_{{\rm kin},I}\lesssim 10$ eV and the cross section corresponds to that of gas probes by increasing $E_{{\rm kin},I}$ atrazhev1981hot ; atrazhev1985heating . Then, for high energy electrons $E_{{\rm kin},I}\gg 10$ eV, the cross section will not so differ from the isolated atom calculation. But, the structure of outermost shell may differ between liquid and isolated probes.
This treatment of the Migdal effect in Ref. Ibe:2017yqa does not include explicit couplings between different electronic channels due to non-adiabatic transitions. The dimensionless order parameter is assumed from the time-dependent perturbation theory as $\kappa_{ab}=|\vec{d}_{ab}\cdot\vec{v}_{n}|/E_{ab}$. Here, $\vec{d}_{ab}$ is the transition matrix of states $a$ and $b$, $\vec{v}_{n}$ is the velocity of nuclear recoil, and $E_{ab}$ is the energy difference. Assuming $|\vec{d}_{ab}\cdot\vec{v}_{n}|\sim a_{0}^{-1}|\vec{v}_{n}|$, the parameter becomes $\kappa_{ab}\sim(7\ {\rm keV}/E_{ab})\sqrt{E_{R}/m_{n}}$. Then, $\kappa_{ab}\ll 1$ will be satisfied at least for $E_{R}\lesssim 1$ keV or $E_{ab}\gtrsim 100$ eV. A fully consistent treatment of the Migdal effect including multi-channel non-adiabatic couplings and time-dependent correlation effects would require a more sophisticated quantum-chemistry or time-dependent multi-electron approach, which will be a future improvement.
Acknowledgements
This work is supported by JSPS Grant-in-Aid for Scientific Research KAKENHI Grant No. JP23KJ2173.